OpenStudy (anonymous):

write an explicit formula for the recursive formula: a1=4a n-1 + 16

OpenStudy (superdavesuper):

hmmm @lashae1995 would u plz check n see if u got everything in the eqn u posted above? seems like its missing something....

OpenStudy (superdavesuper):

Or is it really just a_1 = a_n-1 +16? I think there should be the term a_n somewhere in the equation....

OpenStudy (anonymous):

a1=4, an=a (n-1) +16

OpenStudy (superdavesuper):

Ah thank you :)

OpenStudy (superdavesuper):

a1=4 a2=a1+16 a3=a2+16=a1+16+16 ... a_n=a_(n-1)+16=a1+(n-1)*16 following so far?